OpenStudy (anonymous):

Martha paid $42.40 for her sister’s birthday present. This included 6% sales tax. What was the cost of the gift before tax

OpenStudy (anonymous):

that's not one of answer choices

OpenStudy (anonymous):

somebody plz help me

OpenStudy (anonymous):

Write an equation and solve. Round to the nearest hundredth where necessary. 19 is what percent of 40

OpenStudy (anonymous):

Two U.S. states are not part of the continental United States. What percent of the fifty U.S. states are included in the continental United States

OpenStudy (anonymous):

42.40=x*.06

OpenStudy (anonymous):

42.40 - 6% of 42.40 = original amount before tax 42.40 - (.06)(42.40) = original amount before tax 42.40 - 2.54 = 39.86 check.... 39.96 + 2.54 = 42.40 42.40 = 42.40 (correct) cost before tax was $ 39.86

OpenStudy (anonymous):

19 is what % of 40 19 = a% x 40 19/40 = a% .475 = a% 47.5 = a check... 19 = (.475)(40) 19 = 19 (correct) 19 is 47.5% of 40

OpenStudy (anonymous):

48/50 are continental states .96 = 96% 96% of the states are continental states